A proton makes a (radius = 30.0 cm) circular orbit in a 2.5 T magnetic field. What is the velocity of the proton?

mass of proton = 1.67x10-27 kge = 1.602x10-19 C

1
Expert's answer
2022-05-11T13:53:11-0400

"evB=\frac{mv^2}r,"

"eBr=mv,"

"v=\frac{eBr}m=0.72\cdot 10^8~\frac ms."